1. Cinnamon Buns Co. (CBC) started 2018 with $53,000 of merchandise on hand. During 2018, $285,000 in merchandise was purchased on account with credit terms of 1/10 n/30. All discounts were taken. Purchases were all made f.o.b. shipping point. CBC paid freight charges of $10,700. Merchandise with an invoice amount of $3900 was returned for credit. Cost of goods sold for the year was $313,000. CBC uses a perpetual inventory system.

What is cost of goods available for sale, assuming CBC uses the gross method?

2. Vijay Inc. purchased a three-acre tract of land for a building site for $280,000. On the land was a building with an appraised value of $128,000. The company demolished the old building at a cost of $12,200, but was able to sell scrap from the building for $1690. The cost of title insurance was $860 and attorney fees for reviewing the contract were $430. Property taxes paid were $2100, of which $150 covered the period subsequent to the purchase date. The capitalized cost of the land is:

3. Short Corporation acquired Hathaway, Inc., for $44,050,000. The fair value of all Hathaway's identifiable tangible and intangible assets was $42,000,000. Short will amortize any goodwill over the maximum number of years allowed. What is the annual amortization of goodwill for this acquisition?

4. On March 31, 2018, M. Belotti purchased the right to remove gravel from an old rock quarry. The gravel is to be sold as roadbed for highway construction. The cost of the quarry rights was $174,000, with estimated salable rock of 20,000 tons. During 2018, Belotti loaded and sold 5000 tons of rock and estimated that 15,000 tons remained at December 31, 2018. At January 1, 2019, Belotti estimated that 15,000 tons still remained. During 2019, Belotti loaded and sold 10,000 tons. Belotti uses the units-of-production method.
  
Belotti would record depletion in 2019 of:

5. On July 8, a fire destroyed the entire merchandise inventory on hand of Larrenaga Wholesale Corporation. The following information is available:

Sales, January 1 through July 8 $ 684,000
Inventory, January 1 132,000
Purchases, January 1 through July 8 655,000
Gross profit ratio 28 %

What is the estimated inventory on July 8 immediately prior to the fire?
Answer #1
1)
Beginning inventory $    53,000.00
Inventory purchased $ 285,000.00
Freight $    10,700.00
Merchandise returned $    (3,900.00)
Discounts [($285,000 – 3,900) x 1%)] $    (2,811.00)
Cost of goods available for sale $ 341,989.00
2)
Purchase price    280,000.00
Demolition costs      12,200.00
Scrap Sold      (1,690.00)
Title insurance            860.00
Legal Fees            430.00
Property Taxes        1,950.00
Total cost    293,750.00
3. Zero
4. For 2018 =174,000/20,000 *5,000 = 43,500
For 2019 = (174,000-43,500)/15,000 *10,000 = 87,000
So Belotti would record depletion in 2019 of 87,000
5)
Cost of goods available for sale $ 787,000.00
Cost of goods sold $ 492,480.00
Ending Inventory $ 294,520.00